OpenStudy (anonymous):

Solve for F c=5/9(f-32) enter in this format F=(__)C+__

OpenStudy (texaschic101):

C = 5/9(f - 32) C/(5/9) = f - 32 C/(5/9) + 32 = f I am not sure how to write this...

OpenStudy (texaschic101):

oh..wait.....C/(5/9) = C * 9/5 so I think your answer is : F = 9/5C + 32